# Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

### v2

@@ Acknowledgments
Thanks to Sharat Chikkerur, Mohammad Havaei, and the anonymous reviewers for feedback on this paper. The research also benefited from feedback from David Bruns-Smith, Ming-Wei Chang, Michael Collins, Patrick Fernandez, Mandar Joshi, Rishabh Joshi, Balaji Lakshminarayanan, Kenton Lee, Kristina Toutanova, Victor Veitch, and Zihao Wang. Finally, we thank the people who built the infrastructure used in our experiments, including the T5X team and Léonard Hussenot, Johan Ferret, Robert Dadashi, Geoffrey Cideron, Alexis Jacq, Sabela Ramos, Piotr Stanczyk, Sertan Girgin, Danila Sinopalnikov, Amélie Héliou, Bobak Shahriari, Bilal Piot, Matt Hoffmann, Nikola Momchev, and Olivier Bachem.
@@ Appendix A Numerical results for best-of-n reranking
Average agreement between reward models are shown in Tables 3-6. Autoevaluation results are shown in Table 7 and 8.
@@ Appendix B Numerical results for RLHF
ft median 0.010 2.202 1.480 -1.939
ft median 0.030 1.943 1.596 -0.509
ft median 0.100 0.977 0.809 -0.376
ft median 0.300 -0.249 -0.370 -0.789
ft median 0.500 -0.715 -0.788 -0.963

### v3

@@ Acknowledgments
Thanks to Sharat Chikkerur and Mohammad Havaei for feedback on this paper. The research also benefited from discussions with David Bruns-Smith, Ming-Wei Chang, Michael Collins, Anca Dragan, Chris Dyer, Patrick Fernandez, Mandar Joshi, Rishabh Joshi, Balaji Lakshminarayanan, Kenton Lee, Kristina Toutanova, Victor Veitch, and Zihao Wang. We would also like to acknowledge the reviewers at the Conference on Language Modeling for their input. Finally, we thank the people who built the infrastructure used in our experiments, including the T5X team and Léonard Hussenot, Johan Ferret, Robert Dadashi, Geoffrey Cideron, Alexis Jacq, Sabela Ramos, Piotr Stanczyk, Sertan Girgin, Danila Sinopalnikov, Amélie Héliou, Bobak Shahriari, Bilal Piot, Matt Hoffmann, Nikola Momchev, and Olivier Bachem.
@@ Appendix A Numerical results for best-of-n reranking
Average agreement between reward models are shown in Tables 2-5. Autoevaluation results are shown in Table 6 and 7.
@@ Appendix B Numerical results for RLHF
ft median 0.010 2.206 1.480 -1.939
ft median 0.030 1.939 1.596 -0.509
ft median 0.100 0.955 0.809 -0.376
ft median 0.300 -0.265 -0.370 -0.789
ft median 0.500 -0.728 -0.788 -0.963
@@ Appendix C Hyperparameters
Appendix D Additional results
Task Prompt Output
TL;DR I’ll try to keep this short! **Background** * I’ve always been an on again/off again (very casual!) jogger, typically doing 3 - 5 k * My knees have always been finicky, and I went to a physio who thought I had “runner’s knee” […] I had my baby a year ago, so all in all I haven’t run for about 1.5 years […] However, when I run even the tinyiest bit, or even go on a long walk or a hike, my pelvis gets very sore and tight […] I think pregnancy messed with my body, now I can’t even run even the smallest amount without pain in my pelvis and knees. I’m fairly certain the problem isn’t just that I’m completely out of shape.
Helpfulness Human: Assistant, can you help me find books that have really popular antagonists? Assistant: Sure! I know many books that feature good antagonists, such as those in the Lord of the Rings books. Human: Who specifically from Lord of the Rings? Assistant: Well, Sauron, for example. If you want, I can explain who he was.
XSum/NLI The ex-Reading defender denied fraudulent trading charges relating to the Sodje Sports Foundation - a charity to raise money for Nigerian sport. Mr Sodje, 37, is jointly charged with elder brothers Efe, 44, Bright, 50 and Stephen, 42. Appearing at the Old Bailey earlier, all four denied the offence. The charge relates to offences which allegedly took place between 2008 and 2014. Sam, from Kent, Efe and Bright, of Greater Manchester, and Stephen, from Bexley, are due to stand trial in July. They were all released on bail. Former Premier League footballer Sam Sodje has appeared in court alongside three brothers accused of charity fraud.
(a) tl;dr
(b) helpfulness